# Temporal Valued Constraint Satisfaction Problems

### v1

@@ Abstract.
We study the complexity of the valued constraint satisfaction problem (VCSP) for every valued structure with the domain ${\mathbb{Q}}$ that is preserved by all order-preserving bijections. Such VCSPs will be called temporal, in analogy to the (classical) constraint satisfaction problem: a relational structure is preserved by all order-preserving bijections if and only if all its relations have a first-order definition in $({\mathbb{Q}};<)$, and the CSPs for such structures are called temporal CSPs. Many optimization problems that have been studied intensively in the literature can be phrased as a temporal VCSP. We prove that a temporal VCSP is in P, or NP-complete. Our analysis uses the concept of fractional polymorphisms; this is the first dichotomy result for VCSPs over infinite domains which is complete in the sense that it treats all valued structures with a given automorphism group.
1. Introduction
@@
Valued constraint satisfaction problems (VCSPs) form a large class of computational optimization problems. A VCSP is parameterized by a valued structure (sometimes called the template), which consists of a domain $D$ and cost functions, each defined on $D^{k}$ for some $k$. The input to the VCSP consists of a finite set of variables, a finite sum of cost functions applied to these variables, and a threshold $u$, and the task is to find an assignment to the variables so that the sum of the costs is at most $u$. The computational complexity of such problems has been studied depending on the valued structure that parameterizes the problem. VCSPs generalize constraint satisfaction problems (CSPs), which can be viewed as a variant of VCSPs with costs from the set $\{0,\infty\}$: every constraint is either satisfied or surpasses every finite threshold. VCSPs also generalize min-CSPs, which are the natural variant of CSPs where, instead of asking whether all constraints can be satisfied at once, we search for an assignment that minimizes the number of unsatisfied constraints. Such problems can be modeled as VCSPs with costs from the set $\{0,1\}$.
@@
A major achievement of the field is that if the domain of the valued structure ${\mathfrak{A}}$ is finite, then the computational complexity of $\operatorname{VCSP}({\mathfrak{A}})$ is in P, or NP-complete. This result has an interesting history. The classification task was first considered in [21] with important first results that indicated that we might expect a good systematic theory for such VCSPs. A milestone was reached by Thapper and Živný with the proof of a complexity dichotomy for the case where the cost functions never take value $\infty$ [38]. On the hardness side, Kozik and Ochremiak [29] formulated a condition that implies hardness for $\operatorname{VCSP}({\mathfrak{A}})$ and found equivalent characterisations that suggested that this condition characterises NP-hardness (unless P=NP, of course). Kolmogorov, Krokhin, and Rolínek [28] then showed that if the hardness condition from [29] does not apply, linear programming relaxation in combination with algorithms for classical CSPs can be used to solve $\operatorname{VCSP}({\mathfrak{A}})$, conditional on the tractability conjecture for (classical) CSPs. Finally, this conjecture about CSPs has been confirmed [41] (the result has been announced independently by Bulatov [17] and by Zhuk [42]), thus completing the complexity dichotomy for $\operatorname{VCSP}({\mathfrak{A}})$ for finite-domain templates ${\mathfrak{A}}$ as well. A key tool for distinguishing the tractable VCSPs from the NP-hard ones are fractional polymorphisms, which, in some sense, capture the symmetries of the VCSP template.
@@
For VCSPs with infinite templates we cannot hope for general classification results, since this is already out of reach for the special case of CSPs over infinite domains [6]. However, a powerful algebraic machinery was developed to study CSPs of structures with a rich automorphism group, which has led to classification results for many concrete automorphism groups: we list [9, 12, 11, 4, 31] as a representative sample. Some part of this machinery has also been developed for VCSPs in [15], inspired by the concepts from infinite-domain CSPs and finite-domain VCSPs [26, 29]. Nevertheless, no complexity classification such as for the classes of CSPs discussed above was obtained so far.
In this article we provide the first VCSP dichotomy result for a class of templates which consists of all valued structures preserved by some fixed permutation group. Concretely, we prove that the VCSP for every valued structure with the domain $\mathbb{Q}$ that is preserved by all order-preserving bijections is in P or NP-complete. We call such valued structures temporal, in analogy to temporal relational structures, i.e., structures with the domain $\mathbb{Q}$ preserved by all order-preserving bijections – these are precisely the structures with a first-order definition over $(\mathbb{Q};<)$. We also provide disjoint algebraic conditions that characterize the tractable and the NP-complete case, in analogy to similar classifications for classes of (V)CSPs. The result confirms the dichotomy conjecture from [15, Conjecture 9.3] for the special case of temporal valued structures.
Apart from the relevance of temporal CSPs as a test case for understanding VCSPs on infinite domains, they constitute an important class for several reasons:
  • Temporal CSPs encompass many natural optimization problems [^1] It should be mentioned here that these problems come in two flavours: one is where the input is a graph, or more generally a structure; the other, which we adopt here, is that the input consists of a finite sum of cost functions, which in particular allows that the sum contains identical summands. In the setting of Min-CSPs for graphs, this corresponds to considering multigraphs in the input. However, often (but not always, see Theorems 8.6 and 8.7 in [30]) the two variants of the problem have the same complexity (see, e.g., [22, Section 2.3] for relevant techniques in this context). such as Directed Feedback Arc Set, Directed Subset Feedback Arc Set, Edge Multicut (aka Min-Correlation-Clustering with Partial Information), Symmetric Directed Multicut, Steiner Multicut, Disjunctive Multicut, and many more [16, 23, 32, 34, 33].
@@
Our tractability condition for temporal VCSPs is based on fractional polymorphisms, that is, probability distributions on operations with particular properties. Surprisingly, the fractional polymorphisms that appear in this classification are of a very particular shape: there is always a single operation with probability 1. The operations that appear in this context are the same operations that were already essential for classification of temporal CSPs. The hardness condition is based on the notion of expressibility, which generalizes primitive positive definitions, and on the notion of generalized pp-constructions [15], which provide polynomial-time reductions between VCSPs. Interestingly enough, it is not known whether preservation by fractional polymorphisms characterises expressibility in our setting; this is known for valued structures with a finite domain [20, 26]. Our classification proof, however, does not rely on such a characterisation.
1.1. Related Work
@@
VCSPs on infinite domain have been studied in [39, 37, 40]. Nevertheless, the valued structures considered in these articles typically do not have an oligomorphic automorphism group, a property that is essential for applying our techniques for classifying the complexity of VCSPs. The foundations of the theory for VCSPs of valued structures with an oligomorphic automorphism group were laid in [15] with the motivation to study resilience problems from database theory. Concrete subclasses of temporal VCSPs have been studied in the context of min-CSPs from the parameterized complexity perspective: the complexity of equality min-CSPs has been classified in [35] and parameterized complexity of min-CSPs over the Point Algebra has been classified in [33]. The authors mention a classification of first-order generalizations of the Point Algebra as a natural continuation of their research [33, Section 5]. The present paper contains a classification of VCSPs over all temporal structures and thus lays the foundations for classifying the parameterized complexity of min-CSPs for such structures, including algebraic techniques that we expect to be useful in this context as well.
1.2. Outline
@@
The article is organized as follows. Section 2 contains preliminaries on VCSPs in general, with some notation and properties specific to temporal VCSPs. Section 3 contains several new facts about VCSPs that have been used in the classification. Section 4 contains the classification of equality VCSPs, that is, VCSPs of valued structures with an automorphism group equal to the full symmetric group; on the one hand, this serves as a warm-up, on the other hand it is a building block for the general case. Section 5 contains the full classification of temporal VCSPs, which is the main contribution of the paper. We conclude with some promising questions for future research in Section 6.
2. Preliminaries
@@
Let :=${\mathbb{N}}:=\{0,1,2,\dots\}$ be the set of natural numbers. For $k\in\mathbb{N}$ the set $\{1,\dots,k\}$ will be denoted by $[k]$. The set of rational numbers is denoted by $\mathbb{Q}$ and the standard strict linear order of $\mathbb{Q}$ by $<$. We also need an additional value $\infty$; all we need to know about $\infty$ is that
@@
If $A$ is a set and $t\in A^{k}$, then we implicitly assume that $t=(t_{1},\dots,t_{k})$, where $t_{1},\dots,t_{k}\in A$. If $f\colon A^{\ell}\to A$ is an operation on $A$ and $t^{1},\dots,t^{\ell}\in A^{k}$, then we denote
$(f(t^{1}_{1},t^{2}_{1},\dots,t^{\ell}_{1}),\ldots,f(t^{1}_{k},t^{2}_{k},\dots,t^{\ell}_{k}))$
by $f(t^{1},\dots,t^{\ell})$ and say that $f$ is applied componentwise.
2.1. Valued structures
@@
Usual relations will also be called crisp relations. A valued relation $R\in{\mathscr{R}}_{A}^{(k)}$ that only takes values from $\{0,\infty\}$ will be identified with the crisp relation $\{t\in A^{k}\mid R(t)=0\}.$ A valued relation is called essentially crisp if it attains at most one finite value. For $R\in{\mathscr{R}}_{A}^{(k)}$ the feasibility relation of $R$ is defined as
$\operatorname{Feas}(R):=\{t\in A^{k}\mid R(t)<\infty\}.$
For $S\subseteq A^{k}$ and $a,b\in{\mathbb{Q}}\cup\{\infty\}$, we denote by $S_{a}^{b}$ the valued relation such that $S_{a}^{b}(t)=a$ if $t\in S$, and $S_{a}^{b}(t)=b$ otherwise. We often write $S_{0}^{\infty}$ to stress that $S$ is a crisp relation.
Example 2.1.
@@
On the domain $\mathbb{Q}$, the valued relation $(=)_{0}^{\infty}$ denotes the crisp equality relation, while $(<)_{0}^{1}$ denotes the valued relation $(<)_{0}^{1}(x,y)=0$ if $x<y$ and $(<)_{0}^{1}(x,y)=1$ if $x\geq y$. $(\emptyset)_{0}^{\infty}$ is the unary empty relation (where every element of $\mathbb{Q}$ evaluates to $\infty$).
@@
A (relational) signature $\tau$ is a set of relation symbols, each of them equipped with an arity from ${\mathbb{N}}$. A valued $\tau$-structure ${\mathfrak{A}}$ consists of a set $A$, which is also called the domain of ${\mathfrak{A}}$, and a valued relation $R^{{\mathfrak{A}}}\in{\mathscr{R}}_{A}^{(k)}$ for each relation symbol $R\in\tau$ of arity $k$. All valued structures in this article have countable domains. We often write $R$ instead of $R^{{\mathfrak{A}}}$ if the valued structure is clear from the context. When not specified, we assume that the domains of valued structures ${\mathfrak{A}},{\mathfrak{B}},{\mathfrak{C}},\dots$ are denoted $A,B,C,\dots$, respectively. If ${\mathcal{R}}$ is a set of valued relations over a common domain $A$, we write $(A;{\mathcal{R}})$ for a valued structure ${\mathfrak{A}}$ whose relations are precisely the relations from $\mathcal{R}$; we only use this notation if the precise choice of the signature does not matter. A valued $\tau$-structure where all valued relations only take values from $\{0,\infty\}$ may then be viewed as a relational or crisp $\tau$-structure in the classical sense. A valued structure is called essentially crisp if all of its valued relations are essentially crisp. If ${\mathfrak{A}}$ is a valued $\tau$-structure on the domain $A$, then $\operatorname{Feas}({\mathfrak{A}})$ denotes the relational $\tau$-structure ${\mathfrak{A}}^{\prime}$ on the domain $A$ where $R^{{\mathfrak{A}}^{\prime}}=\operatorname{Feas}(R^{\mathfrak{A}})\in\langle{\mathfrak{A}}\rangle$ for every $R\in\tau$. If $\sigma\subseteq\tau$ and ${\mathfrak{A}}^{\prime}$ is a valued $\sigma$-structure such that $R^{{\mathfrak{A}}^{\prime}}=R^{{\mathfrak{A}}}$ for every $R\in\sigma$, then we call ${\mathfrak{A}}^{\prime}$ a reduct of ${\mathfrak{A}}$.
2.2. Valued constraint satisfaction problems

### v2

# A Complexity Dichotomy for Temporal Valued Constraint Satisfaction Problems

@@ Abstract
We study the computational complexity of the valued constraint satisfaction problem (VCSP) for every valued structure over $\mathbb{Q}$} that is preserved by all order-preserving bijections. Such VCSPs will be called temporal, in analogy to the (classical) constraint satisfaction problem: a relational structure is preserved by all order-preserving bijections if and only if all its relations have a first-order definition in $({\mathbb{Q}};<)$, and the CSPs for such structures are called temporal CSPs. Many optimization problems that have been studied intensively in the literature can be phrased as a temporal VCSP. We prove that a temporal VCSP is in P, or NP-complete. Our analysis uses the concept of fractional polymorphisms. This is the first dichotomy result for VCSPs over infinite domains which is complete in the sense that it treats all valued structures that contain a given automorphism group.
keywords:
Constraint Satisfaction Problems , valued CSPs , temporal CSPs , fractional polymorphisms , complexity dichotomy , min CSPs 68Q25 , 08A70
journal: Information and Computation
[…]
[MB]organization=Institute of Algebra, TU Dresden,city=Dresden, postcode=01062, country=Germany
[…]
[EB]organization=Univ Lyon, CNRS, ENS de Lyon, Université Claude Bernard Lyon 1, LIP UMR5668,city=Lyon, postcode=69342, country=France
[…]
[ZS]organization=Institute of Discrete Mathematics and Geometry, TU Wien,addressline=Karlsplatz 13, city=Wien, postcode=1040, country=Austria
1 Introduction
@@
Valued constraint satisfaction problems (VCSPs) form a large class of computational optimization problems. A VCSP is parameterized by a valued structure (called the template), which consists of a domain $D$ and cost functions, each defined on $D^{k}$ for some $k$. The input to the VCSP consists of a finite set of variables, a finite sum of cost functions applied to these variables, and a threshold $u$, and the task is to find an assignment to the variables so that the sum of the costs is at most $u$. The computational complexity of such problems has been studied depending on the valued structure that parameterizes the problem. VCSPs generalize constraint satisfaction problems (CSPs), which can be viewed as a variant of VCSPs with costs from the set $\{0,\infty\}$: every constraint is either satisfied or surpasses every finite threshold. VCSPs also generalize min-CSPs, which are the natural variant of CSPs where, instead of asking whether all constraints can be satisfied at once, we search for an assignment that minimizes the number of unsatisfied constraints. Such problems can be modeled as VCSPs with costs from the set $\{0,1\}$.
@@
A major achievement of the field is that if the domain of the valued structure $\mathfrak{A}$} is finite, then the computational complexity of $\operatorname{VCSP}({\mathfrak{A}})$ is in P, or NP-complete. This result has an interesting history. The classification task was first considered in [24] with important first results that indicated that we might expect a good systematic theory for such VCSPs. A milestone was reached by Thapper and Živný with the proof of a complexity dichotomy for the case where the cost functions never take value $\infty$ [42]. On the hardness side, Kozik and Ochremiak [32] formulated a condition that implies hardness for $\operatorname{VCSP}({\mathfrak{A}})$ and found equivalent characterisations that suggested that this condition characterises NP-hardness (unless P=NP, of course). Kolmogorov, Krokhin, and Rolínek [31] then showed that if the hardness condition from [32] does not apply, linear programming relaxation in combination with algorithms for classical CSPs can be used to solve $\operatorname{VCSP}({\mathfrak{A}})$, conditional on the tractability conjecture for (classical) CSPs. Finally, this conjecture about CSPs has been confirmed independently by Bulatov [19] and by Zhuk [45, 46], thus completing the complexity dichotomy for $\operatorname{VCSP}({\mathfrak{A}})$ for finite-domain templates $\mathfrak{A}$} as well. A key tool for distinguishing the tractable VCSPs from the NP-hard ones are fractional polymorphisms, which, in some sense, capture the symmetries of the VCSP template.
@@
For VCSPs with infinite templates we cannot hope for general classification results, since this is already out of reach for the special case of CSPs over infinite domains [5]. However, a powerful algebraic machinery was developed to study CSPs of structures with a rich automorphism group, which has led to classification results for many concrete automorphism groups: we list [8, 13, 10, 3, 34] as a representative sample. Some part of this machinery has also been developed for VCSPs in [16], inspired by the concepts from infinite-domain CSPs and finite-domain VCSPs [29, 32]. Nevertheless, no complexity classification for VCSPs such as for the classes of CSPs discussed above was obtained so far.
In this article we provide the first VCSP dichotomy result for a class of templates which consists of all valued structures preserved by some fixed permutation group. Concretely, we prove that the VCSP for every valued structure with the domain $\mathbb{Q}$ that is preserved by all order-preserving bijections is in P or NP-complete. We call such valued structures temporal, in analogy to temporal relational structures, i.e., structures with the domain $\mathbb{Q}$ preserved by all order-preserving bijections – these are precisely the structures with a first-order definition over $(\mathbb{Q};<)$. We also provide disjoint algebraic conditions that characterize the tractable and the NP-complete case, in analogy to similar classifications for classes of (V)CSPs. The result confirms the dichotomy conjecture from [16, Conjecture 9.3] for the special case of temporal valued structures. Note that our classification result is incomparable with the result of [9] which shows that that submodular PLH functions form a maximally tractable class of PLH cost functions, because the class of temporal valued structures is a proper subclass of PLH and contains structures with tractable VCSPs that are not submodular.
Apart from the relevance of temporal VCSPs as a test case for understanding VCSPs on infinite domains, they constitute an important class for several reasons:
  • 1.
    Temporal VCSPs encompass many natural optimization problems [^1] It should be mentioned that these problems come in two flavours: one is where the input is a graph, or more generally a structure; the other, which we adopt here, is that the input consists of a finite sum of cost functions, which in particular allows that the sum contains identical summands. In the setting of Min-CSPs for graphs, this corresponds to considering multigraphs in the input. However, often (but not always, see Theorems 8.6 and 8.7 in [33]) the two variants of the problem have the same complexity (see, e.g., [25, Section 2.3] for relevant techniques in this context)., for example, Directed Feedback Arc Set, Directed Subset Feedback Arc Set, Edge Multicut (aka Min-Correlation-Clustering with Partial Information), Symmetric Directed Multicut, Steiner Multicut, Disjunctive Multicut, and many more [18, 26, 36, 38, 35].
  • 2.
@@
  • 3.
    Temporal VCSPs contain interesting polynomial-time tractable cases with some non-trivial interaction between the ‘soft’ (i.e., finite-valued) and ‘crisp’ constraints (i.e., $\{0,\infty\}$-valued) in the input ( Lemma 47).
@@
Our tractability condition for temporal VCSPs is phrased in terms of fractional polymorphisms, that is, probability distributions on operations with particular properties. Surprisingly, the fractional polymorphisms that appear in this classification are of a very particular shape: there is always a single operation with probability 1. It is known that in general, such fractional polymorphisms are not sufficient to capture the border between polynomial-time tractable and NP-hard VCSPs [22], and we therefore present fractional polymorphisms in full generality for consistency with the literature. The concrete operations that appear in this context are the same operations that were already essential for the the classification of temporal CSPs. The hardness condition is based on the notion of pp-expressibility, which generalizes primitive positive definitions, and on the notion of generalized pp-constructions [16], which provide polynomial-time reductions between VCSPs. Interestingly enough, it is not known whether preservation by fractional polymorphisms characterises pp-expressibility in our setting; this is known for valued structures with a finite domain [23, 29]. Our classification proof, however, does not rely on such a characterisation.
1.1 Related work
@@
VCSPs on infinite domains have been studied in [44, 40, 43]. However, the valued structures considered in these articles typically do not have an oligomorphic automorphism group, a property that is essential for applying our techniques for classifying the complexity of VCSPs. The foundations of the theory for VCSPs of valued structures with an oligomorphic automorphism group were laid in [16] with the motivation to study resilience problems from database theory. Concrete subclasses of temporal VCSPs have been studied in the context of min-CSPs from the parameterized complexity perspective: the complexity of equality min-CSPs has been classified in [37] and parameterized complexity of min-CSPs over the Point Algebra has been classified in [35]. The authors mention a classification of first-order generalizations of the Point Algebra as a natural continuation of their research [35, Section 5]. The present paper contains a classification of VCSPs over all temporal structures and thus provides the foundations for classifying the parameterized complexity of min-CSPs for such structures, including algebraic techniques that we expect to be useful in this context as well.
1.2 Outline
@@
The article is organized as follows. Section 2 contains preliminaries on VCSPs in general, with some notation and properties specific to temporal VCSPs. Section 3 contains several new facts about VCSPs that have been used in the classification. Section 4 contains the classification of equality VCSPs, that is, VCSPs of valued structures with an automorphism group equal to the full symmetric group; on the one hand, this serves as a warm-up, on the other hand it is a building block for the general case. Section 5 contains the full classification of temporal VCSPs, which is the main contribution of the paper.
2 Preliminaries
@@
Let ${\mathbb{N}}:=\{0,1,2,\dots\}$ be the set of natural numbers. For $k\in\mathbb{N}$ the set $\{1,\dots,k\}$ will be denoted by $[k]$. The set of rational numbers is denoted by $\mathbb{Q}$ and the standard strict linear order of $\mathbb{Q}$ by <<. We also need an additional value $\infty$; all we need to know about $\infty$ is that
  • 1.
@@
If $A$ is a set, then $\operatorname{Sym}(A)$ denotes the group of all permutations of $A$. If $t\in A^{k}$, then we implicitly assume that $t=(t_{1},\dots,t_{k})$, where $t_{1},\dots,t_{k}\in A$. For an operation $f\colon A^{\ell}\to A$ and $t^{1},\dots,t^{\ell}\in A^{k}$, we use the following notation for applying $f$ componentwise:
$f(t^{1},\dots,t^{\ell}):=(f(t^{1}_{1},t^{2}_{1},\dots,t^{\ell}_{1}),\ldots,f(t^{1}_{k},t^{2}_{k},\dots,t^{\ell}_{k})).$
2.1 Valued structures
@@
Usual relations $R\subseteq A^{k}$ will also be called crisp relations. A valued relation $R\in{\mathscr{R}}_{A}^{(k)}$ that only takes values from $\{0,\infty\}$ will be identified with the crisp relation $\{t\in A^{k}\mid R(t)=0\}.$ A valued relation $R\in{\mathscr{R}}_{A}^{(k)}$ is called essentially crisp if there exists $a\in\mathbb{Q}$ such that $R(t)\in\{a,\infty\}$ for every $t\in A^{k}$. For $R\in{\mathscr{R}}_{A}^{(k)}$ the feasibility relation of $R$ is defined as $\operatorname{Feas}(R):=\{t\in A^{k}\mid R(t)<\infty\}.$ For $S\subseteq A^{k}$ and $a,b\in{\mathbb{Q}}\cup\{\infty\}$, we denote by $S_{a}^{b}$ the valued relation such that $S_{a}^{b}(t)=a$ if $t\in S$, and $S_{a}^{b}(t)=b$ otherwise. We write $S_{0}^{\infty}$ to stress that $S$ is a crisp relation viewed as a valued relation. The unary empty relation on $A$, where every element of $A$ evaluates to $\infty$, is denoted by $\bot$.
Example 1
@@
On the domain $\mathbb{Q}$, the valued relation $(=)_{0}^{\infty}$ denotes the crisp equality relation, while $(<)_{0}^{1}$ denotes the valued relation $(<)_{0}^{1}(x,y)=0$ if $x<y$ and $(<)_{0}^{1}(x,y)=1$ if $x\geq y$.
@@
A (relational) signature $\tau$ is a set of relation symbols, each of them equipped with an arity from $\mathbb{N}$}. A valued $\tau$-structure $\mathfrak{A}$} consists of a set $A$, which is also called the domain of $\mathfrak{A}$}, and a valued relation $R^{{\mathfrak{A}}}\in{\mathscr{R}}_{A}^{(k)}$ for each relation symbol $R\in\tau$ of arity $k$. All valued structures in this article have countable domains. We often write $R$ instead of $R^{{\mathfrak{A}}}$ if the valued structure is clear from the context. When not specified, we assume that the domains of valued structures ${\mathfrak{A}},{\mathfrak{B}},{\mathfrak{C}},\dots$ are denoted $A,B,C,\dots$, respectively. If $\mathcal{R}$} is a set of valued relations over a common domain $A$, we write $(A;{\mathcal{R}})$ for a valued structure $\mathfrak{A}$} whose relations are precisely the relations from $\mathcal{R}$; we only use this notation if the precise choice of the signature does not matter. A valued $\tau$-structure where all valued relations only take values from $\{0,\infty\}$ may then be viewed as a relational or crisp $\tau$-structure in the classical sense. A valued structure is called essentially crisp if all of its valued relations are essentially crisp. If $\mathfrak{A}$} is a valued $\tau$-structure on the domain $A$, then $\operatorname{Feas}({\mathfrak{A}})$ denotes the relational $\tau$-structure $\mathfrak{A}$$^{\prime}$ on the domain $A$ where $R^{{\mathfrak{A}}^{\prime}}=\operatorname{Feas}(R^{\mathfrak{A}})$ for every $R\in\tau$. If $\sigma\subseteq\tau$ and $\mathfrak{A}$$^{\prime}$ is a valued $\sigma$-structure such that $R^{{\mathfrak{A}}^{\prime}}=R^{{\mathfrak{A}}}$ for every $R\in\sigma$, then we call $\mathfrak{A}$$^{\prime}$ a reduct of $\mathfrak{A}$}.
2.2 Valued constraint satisfaction problems